Minedmusician97:

Mrs. Thomas is designing a rectangular pool for her backyard. The width of the pool is 71/2 feet shorter than twice its length. The perimeter of the pool is 75 feet. What is the length of the pool?

jimthompson5910:

L = length W = width 71/2 = 35.5 The width of the pool is 35.5 ft shorter than twice the length We can say W = 2L - 35.5 to help connect the two variables Then you'll use the perimeter of a rectangle formula P = 2(L+W) Plug in W = 2L-35.5 P = 2(L+W) P = 2(L+2L-35.5) P = 2(3L-35.5) P = 6L - 71 Then plug in P = 75 P = 6L - 71 6L - 71 = P 6L - 71 = 75 Solve for L Do you see how to finish up?